Joanne is 3 years older than amy. Three times joanne's age is 15 years more than twice Amy's age. How old are joanne and Amy?

Answer by ikleyn(52803)

.
If Joanne is J years old, then Amy is (J-3) years old.


Further, the condition says


    3J = 15 + 2*(J-3)


    3J = 15 + 2J - 6

     J = 9.


ANSWER.  Joanne is 9 years old;  Amy is 9-3 = 6 years old.


CHECK.   3J = 3*9 = 27;  15 + 2*6 = 15 + 12 = 27.    ! Correct !